Find the area of a parallelogram with vertices (0, 0), (3, 3), (4, 6), and (7, 9)
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

Shoelace Theorem:

0.....0
3.....3
4.....6
7.....9
0.....0

left column to right column: (0*3) + (3*6) + (4*9) + (7*0) = 54
right column to left column: (0*3) + (3*4) + (6*7) + (9*0) = 54

area = 1/2 * |54 - 54| = 0

Answer by ikleyn(52788)

.

It is clear from the first glance that the answer/(the solution) by @Penguin is not correct.

It is seen from the fact that the four given points are not collinear.

So, the solution by @Penguin is wrong.


Where the mistake is ?

--- The mistake is in incorrect using the Shoelace theorem.


To apply this formula, the vertices must be listen in the clockwise (or anti-clockwise) order.

In this case, @Penguin failed to apply the formula correctly - it is the mistake' source.